A Bibliography of The Princeton Companion to Applied Mathematics

Nicholas John Higham School of Mathematics University of Manchester Manchester M13 9PL

UK

Tel: +44 (0)161 275 5800 FAX: +44 (0)161 275 5819

E-mail: nick.higham@manchester.ac.uk (Internet)

12 September 2016 Version 1.01

Abstract

This is a bibliography of publications in the book The Princeton Companion to Applied Mathematics.

Title word cross-reference

N [Saa15a, Yin15]. W [CJ15].

-Body [Saa15a, Yin15].

Adaptation [KR15]. Aided [Tuc15a]. Aircraft [Cha15a]. Airport [Lio15]. Algebra [Bro15, Hig15o]. Algebraic [Sot15]. Algorithms [Hig15b]. Analysis [CS15a, Hig15o, Kam15a, Mar15f, Mor15, ØS15, Tuc15a, Tuc15b]. Applications [Bro15, Bud15, ØS15, Str15a]. Applied [BB15, BGSS15, Hig15a, Hig15j, Hig15m, HDG^{+}15, Hig15q, Hig15u, Ste15b, Win15]. Approach [Gra15]. Approximation [Cuy15]. Arithmetic [Hig15h]. Astronomy [Saa15a]. Asymptotics [Mar15a, Mil15]. Attitudes [Gri15c]. Automatic [Gri15a]. Avatars [KR15].

Baggage [Lio15]. Ball [Arn15]. Bayesian [Hig15a]. Beliefs [Gri15c]. Benford [Hil15]. Bessel [Mar15b]. Bifurcation [Gle15a]. Big [Ach15b]. Biology [Nie15a]. Biomechanics [Jen15]. Black [Hig15c]. Body [Saa15a, Yin15]. Book [Ste15a]. Boundary [Mar15c]. Bubbles [Pro15]. Buildings [BBC15]. Burgers [Mar15d].

Cahn [NC15]. Calculations [Can15]. Calculus [FL15]. Canonical [Hig15l]. Cardiac [Pan15]. Case [Fit15]. Cauchy [Mar15e]. Chains [Mei15a]. Changing [Fit15]. Chaos [Gle15b]. Chemical [Fei15b]. Chinese [xY15]. Chip [HHV15]. Classical [Ton15]. Cloaking [BL15]. Color [Hig15d]. Combinatorial [Vyg15]. Combinatorics [Win15]. Communication [Wil15]. Companion [HDG^{+}15]. Complex [Gle15c, Mar15f]. Compressed [Eld15]. Computation [Tur15]. Computational [Key15a]. Computer [Tuc15a]. Computer-Aided [Tuc15a]. Computing [Don15]. Conformal [Cro15a]. Conservation [Den15b, Key15b]. Continuous [Wri15]. Continuum [Jam15]. Control [BBC15, Hig15e, RÅ15]. Convexity [Hen15]. Convolution [HM15]. Cosmology [Ell15]. Counterterrorism [Gri15c]. Culture [Men15].

Dark [Saa15a]. Data [Kam15a]. Decomposition [Hig15r]. Delta [Upt15]. Design [HHV15]. Differences [Hig15f]. Differential [Eva15, HL15, Mei15b, Sül15]. Differentiation [Gri15a]. Diffusion [Mar15g]. Digital [Hig15d]. Dimensional [CS15a]. Dirac [Den15a]. Diseases [BvdD15]. Divergent [BH15]. Dynamical [Hol15]. Dynamics [Bud15, Mof15a, Shu15].

Earth [Shu15, Sni15]. Economics [Eke15]. Effective [McP15]. Efficient [BBC15]. Einstein [Mac15]. Electronic [Can15]. Element [Hig15g]. Elementary [dDM15]. Energy [BBC15]. Energy-Efficient [BBC15]. Environment [Fit15]. Equation [Ber15, Cha15c, Che15, Den15a, Den15d, Gle15f, Her15, Hig15c, Lau15, Mar15d, Mar15g, Mar15k, Mar15n, NC15]. Equations [Cla15, Den15c, Eva15, Gle15d, Gle15g, HL15, Hig15s, Kre15, Mac15, Mar15e, Mar15h, Mar15l, Mei15b, Mof15b, Ray15, Sül15]. Ergodicity [Gle15b]. Euler [Gle15d, Mar15h]. Europe [Est15]. Evaluating [dDM15]. Evolving [Gri15c]. Experience [Est15]. Experimental [BB15].

Fast [Roc15, Yin15]. Field [Mac15]. Film [Ber15]. Financial [CS15b]. Finite [Hig15f, Hig15g]. Finite-Element [Hig15g]. Flame [Mat15]. Flight [Arn15, Wan15]. Floating [Hig15h]. Floating-Point [Hig15h]. Flows [God15]. Fluid [Mof15a]. Foams [WH15]. Form [Hig15l]. Formation [Sch15]. Fourier [Roc15]. Framework [Str15a]. France [Est15]. Function [CJ15, Fei15a, Mar15i, Upt15]. Functions [GV15, Hig15i, Mar15b, Tem15, Upt15, dDM15].

Gamma [Mar15i]. General [Ell15, Ste15a]. Generalized [Upt15]. Generation [Nie15b]. Geometry [Gra15, Sot15]. Ginzburg [Cha15c]. Goals [Hig15j]. Golf [Arn15]. Government [Est15, xY15]. Granular [God15]. Graph [DH15, Dav15b, Win15]. Green’s [Sni15].

High [Don15]. High-Performance [Don15]. Hilliard [NC15]. History [BGSS15]. Homogenization [San15a]. Hooke [Mar15j]. Hybrid [Gle15e, Gra15].

Ice [Gol15]. Image [Sap15]. Imaging [CB15, Eps15, Hig15d, Sni15]. Impact [Bud15]. Industrial [Lev15b]. Inerters [Smi15]. Infectious [BvdD15]. Inference [Hig15a]. Influence [Est15, xY15]. Information [Ver15]. Insect [Wan15]. Inspires [Lev15b]. Integral [HM15, Kre15]. Interest [Ste15a]. Interval [Tuc15a, Tuc15b]. Invariants [Den15b]. Inverse [SS15]. Inverted [Ach15a].

Jordan [Hig15l].

Kinetic [VM15]. Kit [McK15]. Knotting [Buc15]. Korteweg [Her15]. Krylov [Sim15].

Lagrange [Gle15d]. Lambert [CJ15].

Landau [Cha15c]. Language [Hig15m]. Languages [Hig15q]. Laplace [Mar15k]. Law [Hil15, Mar15j]. Laws [Den15b, Key15b]. Layer [Mar15c]. Level [San15b]. Linear [Hig15o, Wri15]. Linking [Buc15]. Location [BBC15]. Logistic [Gle15f]. Lorenz [Gle15g]. Lyapunov [Hig15s].

Macromolecules [Buc15]. Magnetohydrodynamics [Hug15]. Making [Fit15]. Manifolds [Den15e]. Many [Str15a]. Mapping [Cro15a]. Markov [Mei15a]. Mathematical [DS15, Eke15, Erm15, Gow15, Hig15j, Jen15, Lay15, Lev15b, Sap15]. Mathematician [Est15]. Mathematicians [xY15]. Mathematics [BB15, BGSS15, CS15b, Fit15, Gol15, Hig15a, Hig15m, HDG^{+}15, Hig15q, Hig15u, KR15, Lev15b, Men15, Ste15a, Ste15b]. Mathieu [GV15]. Matrices [Hig15i]. Matrix [Hig15o, Kea15]. Matter [Kam15b, Men15, Saa15a]. Max [Bro15]. Max-Plus [Bro15]. Maxwell [Den15c]. Mechanics [Fre15, Gri15b, Jam15, Ton15]. Mediated [Men15]. Medical [Eps15]. Medium [McP15]. Method [Hig15g, Ray15, San15b, Yin15]. Methods [Hig15n]. Mining [Kam15a, MB15]. Model [Ben15]. Modeling [Gra15, LeV15a, Lev15b, McK15, Pan15, San15c, Tur15]. Modern [Alo15]. Multipole [Yin15]. Multiscale [San15c].

Navier [Mof15b]. Network [Mor15]. Networks [Gri15c, Wil15]. Neuroscience [Erm15]. Newton’s [Ray15]. Noise [Cha15a]. Nonlinear [Ray15, Wri15]. Nonsmooth [Bud15]. Number [Nie15b]. Numeric [Gra15]. Numerical [HL15, Haw15, Hig15o, Lyn15, Sül15].

Optics [Alo15]. Optimal [BBC15].

Optimization [Vyg15, Wri15]. Ordinary [HL15, Mei15b]. Orthogonal [Hig15p].

Pages [GC15]. Painlevé [Cla15]. Paper [Hig15k]. Partial [Eva15, Sül15]. Pattern [Sch15]. Pendulums [Ach15a]. Performance [Don15]. Personal [Est15]. Perturbation [Mil15]. Phenomena [Mar15o]. Physics [Can15]. Physiology [Lay15]. Picture [Ach15b]. Plus [Bro15]. Point [Hig15h]. Policy [Cro15b]. Polynomials [Hig15p]. Popular [Men15]. Portfolio [BLN15]. Prediction [Lyn15]. Pregnancy [McK15]. Princeton [HDG^{+}15]. Principle [Hig15t]. Problem [Coo15, Saa15a, Yin15]. Problems [SS15]. Processing [Gra15, MP15, Sap15]. Programming [Hig15q, Wri15]. Projects [Tur15]. Proofs [Tuc15a]. Propagation [Mat15].

Quantification [MW15]. Quantum

[Gri15b].

Radar [CB15]. Random [Kea15, Nie15b]. Random-Matrix [Kea15]. Ranking [GC15]. Rapidly [Fit15]. Rattling [Bud15]. Ray [Lio15]. Reactions [Fei15b]. Read [Hig15k]. Reduction [Ben15]. Relativity [Ell15, Haw15]. Representations [Men15]. Reproducible [DS15]. Research [DS15, Fit15, Hig15j]. Riccati [Lau15]. Riemann [Mar15e]. Robotics [Wam15].

Salesman [Coo15]. Scaling [CS15a]. Scholes [Hig15c]. Schrödinger [Den15d]. Science [Cro15b, Key15a]. Sciences [DS15]. Screening [Lio15]. Sea [Gol15]. Searching [Dav15b]. Sensing [Eld15]. Sensor [BBC15]. Series [BH15]. Set [San15b]. Shallow [Mar15l]. Shallow-Water [Mar15l]. Shock [Cha15b]. Shocks [Key15c]. SIAM [Cro15b]. Signal [MP15]. Singular [Hig15r]. Singularities [Mar15m]. Sliding [Bud15]. Slipping [Bud15]. Social [Gri15c]. Soft [Kam15b]. Solid [Can15]. Solids [Fre15]. Solution [HL15, Hig15n, Sül15]. Spaces [Fei15a, Hig15d]. Special [Tem15]. Spectral [Dav15a]. Sport [PQ15]. Spread [BvdD15]. State [Can15]. States [Cro15b]. Stochastic [ØS15]. Stokes [Mof15b]. Structure [Can15]. Subspaces [Sim15]. Sylvester [Hig15s]. Symbolic [Gra15]. Symmetric [Str15a]. Symmetry [Ste15b]. System [Shu15]. Systems [Gle15c, Gle15e, Hol15, Nie15a, Saa15b].

Tail [BH15]. Taming [BH15]. Tasks [Lev15b]. Teach [Str15b]. Ten [KR15]. Tensors [Den15e]. Testing [McK15]. Text [MB15]. Theorem [Sni15]. Theories [McP15]. Theory [BLN15, Cuy15, Dav15a, DH15, Gle15a, Kea15, Mil15, RÅ15, Ver15, VM15, Win15]. These [Men15]. Thin [Ber15]. Thin-Film [Ber15]. Tomography [Lio15]. Transform [Roc15]. Transforms [HM15]. Traveling [Coo15]. Tricomi [Che15]. Tsunami [LeV15a]. Turbulence [Hun15].

U.K. [Fit15]. Uncertainty [MW15]. Understand [Hig15k]. United [Cro15b]. Using [Sni15].

Value [Hig15r]. Variational [Hig15t]. Variations [FL15]. via [Tuc15a]. View [Hig15q]. Vishnu [KR15]. Visualization [PJ15]. Voting [Saa15b]. Vries [Her15].

Water [Mar15l]. Wave [Mar15n, Mar15o]. Waves [Cha15b]. Weather [Lyn15]. Web [GC15]. Workflow [Hig15v]. Write [Ste15a]. Writing [Gow15].

X [Lio15]. X-Ray [Lio15].

References

Acheson:2015:IP

[Ach15a] David Acheson. Inverted pendulums. In Higham et al. [HDG^{+}15], pages 741–743. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                                  .P75 2015. URL http://press.
                                                  princeton.edu/titles/10592.
                                                  html.

Acheson:2015:WBP

[Ach15b] David Acheson. What’s the big picture? In Higham et al. [HDG^{+}15], pages 933–935. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                                  /press.princeton.edu/titles/
                                                  10592.html.

Alonso:2015:MO

[Alo15] Miguel A. Alonso. Modern optics. In Higham et al. [HDG^{+}15], pages 673–680. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                                  .P75 2015. URL http://press.
                                                  princeton.edu/titles/10592.
                                                  html.

Arnold:2015:FGB

[Arn15] Douglas N. Arnold. The flight of a golf ball. In Higham et al. [HDG^{+}15], pages 746– 749. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                                  URL http://press.princeton.
                                                  edu/titles/10592.html.

Bailey:2015:EAM

[BB15] David H. Bailey and Jonathan M. Borwein. Experimental applied mathematics. In Higham et al. [HDG^{+}15], pages 925–933. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Borggaard:2015:OSL

[BBC15] Jeffrey T. Borggaard, John A.

Burns, and Eugene M. Cliff. Optimal sensor location in the control of energy-efficient buildings. In Higham et al. [HDG^{+}15], pages 763–767. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Benner:2015:MR

[Ben15] Peter Benner. Model reduction. In Higham et al. [HDG^{+}15], pages 117–119. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Bernoff:2015:TFE

[Ber15] Andrew J. Bernoff. The thinfilm equation. In Higham et al. [HDG^{+}15], pages 169–170. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Barrow-Green:2015:HAM

[BGSS15] June Barrow-Green and Reinhard Siegmund-Schultze. The history of applied mathematics. In Higham et al. [HDG^{+}15], pages 55–79. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Berry:2015:DST

[BH15] Michael V. Berry and Christopher J. Howls. Divergent series: Taming the tail. In Higham et al. [HDG^{+}15], pages 634– 640. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Bryan:2015:C

[BL15] Kurt Bryan and Tanya Leise. Cloaking. In Higham et al. [HDG^{+}15], pages 733–735. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Brennan:2015:PT

[BLN15] Thomas J. Brennan, Andrew W. Lo, and Tri-Dung Nguyen. Portfolio theory. In Higham et al. [HDG^{+}15], pages 648–658. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Broomhead:2015:AMP

[Bro15] David S. Broomhead. Applications of max-plus algebra. In Higham et al. [HDG^{+}15], pages 795– 800. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Buck:2015:KLM

[Buc15] Dorothy Buck. Knotting and linking of macromolecules. In Higham et al. [HDG^{+}15], pages 752– 755. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Budd:2015:SSR

[Bud15] Chris Budd. Slipping, sliding, rattling, and impact: Nonsmooth dynamics and its applications. In Higham et al. [HDG^{+}15], pages 769–771. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Brauer:2015:SID

[BvdD15] Fred Brauer and P. van den Driessche. The spread of infectious diseases. In Higham et al. [HDG^{+}15], pages 687– 694. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Cances:2015:ESC

[Can15] Eric Cancès. Electronic structure calculations (solid state physics). In Higham et al. [HDG^{+}15], pages 847–851. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Cheney:2015:RI

[CB15] Margaret Cheney and Brett Borden. Radar imaging. In Higham et al. [HDG^{+}15], pages 860– 864. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Chapman:2015:AN

[Cha15a] C. J. Chapman. Aircraft noise. In Higham et al. [HDG^{+}15], pages 783–786. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Chapman:2015:SW

[Cha15b] C. J. Chapman. Shock waves. In Higham et al. [HDG^{+}15], pages 720–724. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Chapman:2015:GLE

[Cha15c] S. Jonathan Chapman. The Ginzburg–Landau equation. In Higham et al. [HDG^{+}15], pages 148–149. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Chen:2015:TE

[Che15] Gui-Qiang G. Chen. The Tricomi equation. In Higham et al. [HDG^{+}15], pages 170–171. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Corless:2015:LF

[CJ15] Robert M. Corless and David J. Jeffrey. The Lambert W function. In Higham et al. [HDG^{+}15], pages 151–155. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Clarkson:2015:PE

[Cla15] Peter A. Clarkson. The Painlevé equations. In Higham et al. [HDG^{+}15], pages 163–165. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Cook:2015:TSP

[Coo15] William Cook. The traveling salesman problem. In Higham et al. [HDG^{+}15], pages 778– 781. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Crowdy:2015:CM

[Cro15a] Darren Crowdy. Conformal mapping. In Higham et al. [HDG^{+}15], pages 84–86. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Crowley:2015:SSP

[Cro15b] Jim Crowley. SIAM and science policy in the United States. In Higham et al. [HDG^{+}15], pages 956–959. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Calvetti:2015:DAS

[CS15a] Daniela Calvetti and Erkki Somersalo. Dimensional analysis and scaling. In Higham et al. [HDG^{+}15], pages 90–93. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Carmona:2015:FM

[CS15b] René Carmona and Ronnie Sircar. Financial mathematics. In Higham et al. [HDG^{+}15], pages 640– 648. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Cuyt:2015:AT

[Cuy15] Annie Cuyt. Approximation theory. In Higham et al. [HDG^{+}15], pages 248–262. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Davies:2015:ST

[Dav15a] E. Brian Davies. Spectral theory. In Higham et al. [HDG^{+}15], pages 236–248. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Davis:2015:SG

[Dav15b] Timothy A. Davis. Searching a graph. In Higham et al. [HDG^{+}15], pages 757–759. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

deDinechin:2015:EEF

[dDM15] Florent de Dinechin and Jean-Michel Muller. Evaluating elementary functions. In Higham et al. [HDG^{+}15], pages 759– 761. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Dennis:2015:DE

[Den15a] Mark R. Dennis. The Dirac equation. In Higham et al. [HDG^{+}15], pages 142–144. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Dennis:2015:ICL

[Den15b] Mark R. Dennis. Invariants and conservation laws. In Higham et al. [HDG^{+}15], pages 106– 112. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Dennis:2015:ME

[Den15c] Mark R. Dennis. Maxwell’s equations. In Higham et al. [HDG^{+}15], pages 160–162. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Dennis:2015:SE

[Den15d] Mark R. Dennis. Schrödinger’s equation. In Higham et al. [HDG^{+}15], page 167. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Dennis:2015:TM

[Den15e] Mark R. Dennis. Tensors and manifolds. In Higham et al. [HDG^{+}15], pages 127–130. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Davis:2015:GT

[DH15] Timothy A. Davis and Yifan Hu. Graph theory. In Higham et al. [HDG^{+}15], pages 101– 103. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Dongarra:2015:HPC

[Don15] Jack Dongarra. High-performance computing. In Higham et al. [HDG^{+}15], pages 839–843. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Donoho:2015:RRM

[DS15] David L. Donoho and Victoria Stodden. Reproducible research in the mathematical sciences. In Higham et al. [HDG^{+}15], pages 916–925. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Ekeland:2015:ME

[Eke15] Ivar Ekeland. Mathematical economics. In Higham et al. [HDG^{+}15], pages 868–873. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Eldar:2015:CS

[Eld15] Yonina C. Eldar. Compressed sensing. In Higham et al. [HDG^{+}15], pages 823–827. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Ellis:2015:GRC

[Ell15] George F. R. Ellis. General relativity and cosmology. In Higham et al. [HDG^{+}15], pages 579– 590. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Epstein:2015:MI

[Eps15] Charles L. Epstein. Medical imaging. In Higham et al. [HDG^{+}15], pages 816–823. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Ermentrout:2015:MN

[Erm15] Bard Ermentrout. Mathematical neuroscience. In Higham et al. [HDG^{+}15], pages 873– 879. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Esteban:2015:PEF

[Est15] Maria Esteban. A personal experience in France and Europe of how to influence government as a mathematician. In Higham et al. [HDG^{+}15], pages 954– 956. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Evans:2015:PDE

[Eva15] Lawrence C. Evans. Partial differential equations. In Higham et al. [HDG^{+}15], pages 190– 200. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Feichtinger:2015:FS

[Fei15a] Hans G. Feichtinger. Function spaces. In Higham et al. [HDG^{+}15], pages 99–101. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Feinberg:2015:CR

[Fei15b] Martin Feinberg. Chemical reactions. In Higham et al. [HDG^{+}15], pages 627–634. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Fitt:2015:MCU

[Fit15] Alistair D. Fitt. Making the case for U.K. mathematics research in a rapidly changing environment. In Higham et al.

[HDG^{+}15], pages 959–961. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Fonseca:2015:CV

[FL15] Irene Fonseca and Giovanni Leoni.

Calculus of variations. In Higham et al. [HDG^{+}15], pages 218– 226. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Freund:2015:MS

[Fre15] L. B. Freund. Mechanics of solids. In Higham et al. [HDG^{+}15], pages 505–516. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Gleich:2015:RWP

[GC15] David F. Gleich and Paul G. Constantine. Ranking web pages. In Higham et al. [HDG^{+}15], pages 755–757. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Glendinning:2015:BT

[Gle15a] Paul Glendinning. Bifurcation theory. In Higham et al. [HDG^{+}15], pages 393–402. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Glendinning:2015:CE

[Gle15b] Paul Glendinning. Chaos and ergodicity. In Higham et al. [HDG^{+}15], pages 82–83. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Glendinning:2015:CS

[Gle15c] Paul Glendinning. Complex systems. In Higham et al. [HDG^{+}15], pages 83–84. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Glendinning:2015:ELE

[Gle15d] Paul Glendinning. The Euler– Lagrange equations. In Higham et al. [HDG^{+}15], page 147. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Glendinning:2015:HS

[Gle15e] Paul Glendinning. Hybrid systems. In Higham et al. [HDG^{+}15], pages 103–104. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Glendinning:2015:LEa

[Gle15f] Paul Glendinning. The logistic equation. In Higham et al. [HDG^{+}15], pages 156–158. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Glendinning:2015:LEb

[Gle15g] Paul Glendinning. The Lorenz equations. In Higham et al. [HDG^{+}15], pages 158–159. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Goddard:2015:GF

[God15] Joe D. Goddard. Granular flows.

In Higham et al. [HDG^{+}15], pages 665–673. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Golden:2015:MSI

[Gol15] Kenneth M. Golden. The mathematics of sea ice. In Higham et al. [HDG^{+}15], pages 694– 705. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Gowers:2015:MW

[Gow15] Timothy Gowers. Mathematical writing. In Higham et al. [HDG^{+}15], pages 897–903. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Grandine:2015:HSN

[Gra15] Thomas A. Grandine. A hybrid symbolic–numeric approach to geometry processing and modeling. In Higham et al. [HDG^{+}15], pages 787–790. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Griewank:2015:AD

[Gri15a] Andreas Griewank. Automatic differentiation. In Higham et al. [HDG^{+}15], pages 749–752. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Griffiths:2015:QM

[Gri15b] David Griffiths. Quantum mechanics. In Higham et al. [HDG^{+}15], pages 411–419. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Grindrod:2015:ESN

[Gri15c] Peter Grindrod. Evolving social networks, attitudes, and beliefs— and counterterrorism. In Higham et al. [HDG^{+}15], pages 800– 804. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Gutierrez-Vega:2015:MF

[GV15] Julio C. Gutiérrez-Vega. Mathieu functions. In Higham et al. [HDG^{+}15], pages 159–160. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Hawke:2015:NR

[Haw15] Ian Hawke. Numerical relativity. In Higham et al. [HDG^{+}15], pages 680–687. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Higham:2015:PCA

[HDG^{+}15] Nicholas J. Higham, Mark R. Dennis, Paul Glendinning, Paul A. Martin, Fadil Santosa, and Jared Tanner, editors. The Princeton Companion to Applied Mathematics. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-15039-7 (hardcover). xvii + 3 + 994 + 16 pp. LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Henrion:2015:C

[Hen15] Didier Henrion. Convexity. In Higham et al. [HDG^{+}15], pages 89–90. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Hereman:2015:KVE

[Her15] Willy A. Hereman. The Korteweg– de Vries equation. In Higham et al. [HDG^{+}15], pages 150– 151. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Held:2015:CD

[HHV15] Stephan Held, Stefan Hougardy, and Jens Vygen. Chip design. In Higham et al. [HDG^{+}15], pages 804–808. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Higham:2015:BIA

[Hig15a] Desmond J. Higham. Bayesian inference in applied mathematics. In Higham et al. [HDG^{+}15], pages 658–661. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Higham:2015:A

[Hig15b] Nicholas J. Higham. Algorithms. In Higham et al. [HDG^{+}15], pages 40–49. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Higham:2015:BSE

[Hig15c] Nicholas J. Higham. The Black– Scholes equation. In Higham et al. [HDG^{+}15], pages 137– 138. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Higham:2015:CSD

[Hig15d] Nicholas J. Higham. Color spaces and digital imaging. In Higham et al. [HDG^{+}15], pages 808– 813. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Higham:2015:C

[Hig15e] Nicholas J. Higham. Control. In Higham et al. [HDG^{+}15], pages 88–89. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Higham:2015:FD

[Hig15f] Nicholas J. Higham. Finite differences. In Higham et al. [HDG^{+}15], pages 95–96. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Higham:2015:FEM

[Hig15g] Nicholas J. Higham. The finiteelement method. In Higham et al. [HDG^{+}15], page 96. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Higham:2015:FPA

[Hig15h] Nicholas J. Higham. Floatingpoint arithmetic. In Higham et al. [HDG^{+}15], pages 96–97. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Higham:2015:FM

[Hig15i] Nicholas J. Higham. Functions of matrices. In Higham et al. [HDG^{+}15], pages 97–99. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Higham:2015:GAM

[Hig15j] Nicholas J. Higham. Goals of applied mathematical research. In Higham et al. [HDG^{+}15], pages 48–55. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Higham:2015:HRU

[Hig15k] Nicholas J. Higham. How to read and understand a paper. In Higham et al. [HDG^{+}15], pages 903–906. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Higham:2015:JCF

[Hig15l] Nicholas J. Higham. The Jordan canonical form. In Higham et al. [HDG^{+}15], pages 112– 113. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Higham:2015:LAM

[Hig15m] Nicholas J. Higham. The language of applied mathematics. In Higham et al. [HDG^{+}15], pages 8–27. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Higham:2015:MS

[Hig15n] Nicholas J. Higham. Methods of solution. In Higham et al. [HDG^{+}15], pages 27–40. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Higham:2015:NLA

[Hig15o] Nicholas J. Higham. Numerical linear algebra and matrix analysis. In Higham et al. [HDG^{+}15], pages 263–281. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Higham:2015:OP

[Hig15p] Nicholas J. Higham. Orthogonal polynomials. In Higham et al. [HDG^{+}15], page 122. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Higham:2015:PLA

[Hig15q] Nicholas J. Higham. Programming languages: An applied mathematics view. In Higham et al. [HDG^{+}15], pages 828–839. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Higham:2015:SVD

[Hig15r] Nicholas J. Higham. The singular value decomposition. In Higham et al. [HDG^{+}15], pages 126– 127. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Higham:2015:SLE

[Hig15s] Nicholas J. Higham. The Sylvester and Lyapunov equations. In Higham et al. [HDG^{+}15], pages 168–169. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Higham:2015:VP

[Hig15t] Nicholas J. Higham. Variational principle. In Higham et al. [HDG^{+}15], page 134. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Higham:2015:WAM

[Hig15u] Nicholas J. Higham. What is applied mathematics? In Higham et al. [HDG^{+}15], pages 1–8. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Higham:2015:W

[Hig15v] Nicholas J. Higham. Workflow. In Higham et al. [HDG^{+}15], pages 912–916. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Hill:2015:BL

[Hil15]  Theodore P. Hill.  Benford’s

law. In Higham et al. [HDG^{+}15], pages 135–137. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Hairer:2015:NSO

[HL15] Ernst Hairer and Christian Lubich. Numerical solution of ordinary differential equations. In Higham et al. [HDG^{+}15], pages 293–305. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Higham:2015:ITC

[HM15] Nicholas J. Higham and P. A.

Martin. Integral transforms and convolution. In Higham et al. [HDG^{+}15], pages 104–105. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Holmes:2015:DS

[Hol15] Philip Holmes. Dynamical systems. In Higham et al. [HDG^{+}15], pages 383–393. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Hughes:2015:M

[Hug15] David W. Hughes. Magnetohydrodynamics. In Higham et al. [HDG^{+}15], pages 476– 485. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Hunt:2015:T

[Hun15] Julian C. R. Hunt. Turbulence. In Higham et al. [HDG^{+}15], pages 724–732. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

James:2015:CM

[Jam15] Richard D. James. Continuum mechanics. In Higham et al. [HDG^{+}15], pages 446– 458. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Jensen:2015:MB

[Jen15] Oliver E. Jensen. Mathematical biomechanics. In Higham et al. [HDG^{+}15], pages 609– 616. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Kamath:2015:DMA

[Kam15a] Chandrika Kamath. Data mining and analysis. In Higham et al. [HDG^{+}15], pages 350– 360. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Kamien:2015:SM

[Kam15b] Randall D. Kamien. Soft matter. In Higham et al. [HDG^{+}15], pages 516–523. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Keating:2015:RMT

[Kea15] Jonathan Peter Keating. Randommatrix theory. In Higham et al. [HDG^{+}15], pages 419–428. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Keyes:2015:CS

[Key15a] David E. Keyes. Computational science. In Higham et al. [HDG^{+}15], pages 335–350. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Keyfitz:2015:CL

[Key15b] Barbara Lee Keyfitz. Conservation laws. In Higham et al. [HDG^{+}15], pages 86–88. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Keyfitz:2015:S

[Key15c] Barbara Lee Keyfitz. Shocks. In Higham et al. [HDG^{+}15], pages 122–124. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Krakauer:2015:MAT

[KR15] David Krakauer and Daniel N.

Rockmore. The mathematics of adaptation (or the ten avatars of Vishnu). In Higham et al. [HDG^{+}15], pages 591–597. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Kress:2015:IE

[Kre15] Rainer Kress. Integral equations. In Higham et al. [HDG^{+}15], pages 200–208. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Laub:2015:RE

[Lau15] Alan J. Laub. The Riccati equation. In Higham et al. [HDG^{+}15], pages 165–167. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Layton:2015:MP

[Lay15] Anita T. Layton. Mathematical physiology. In Higham et al. [HDG^{+}15], pages 616–623. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

LeVeque:2015:TM

[LeV15a] Randall J. LeVeque. Tsunami modeling. In Higham et al. [HDG^{+}15], pages 712–720. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Levy:2015:IMI

[Lev15b] Rachel Levy. Industrial mathematics inspires mathematical modeling tasks. In Higham et al.

[HDG^{+}15], pages 940–943. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Lionheart:2015:ABS

[Lio15] W. R. B. Lionheart. Airport baggage screening with X-ray tomography. In Higham et al. [HDG^{+}15], pages 866–868. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Lynch:2015:NWP

[Lyn15] Peter Lynch. Numerical weather prediction. In Higham et al. [HDG^{+}15], pages 705–712. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

MacCallum:2015:EFE

[Mac15] Malcolm A. H. MacCallum. Einstein’s field equations. In Higham et al. [HDG^{+}15], pages 144– 146. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Martin:2015:A

[Mar15a] P. A. Martin. Asymptotics. In Higham et al. [HDG^{+}15], pages 81–82. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Martin:2015:BF

[Mar15b] P. A. Martin. Bessel functions. In Higham et al. [HDG^{+}15], page 137. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Martin:2015:BL

[Mar15c] P. A. Martin. Boundary layer. In Higham et al. [HDG^{+}15], page 82. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL

                                        http://press.princeton.edu/
                                        titles/10592.html.

Martin:2015:BE

[Mar15d] P. A. Martin. The Burgers equation. In Higham et al. [HDG^{+}15], page 138. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Martin:2015:CRE

[Mar15e] P. A. Martin. The Cauchy– Riemann equations. In Higham et al. [HDG^{+}15], page 139. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Martin:2015:CA

[Mar15f] P. A. Martin. Complex analysis. In Higham et al. [HDG^{+}15], pages 173–181. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Martin:2015:DE

[Mar15g] P. A. Martin. The diffusion equation. In Higham et al. [HDG^{+}15], page 142. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Martin:2015:EE

[Mar15h] P. A. Martin. The Euler equations. In Higham et al. [HDG^{+}15], pages 146–147. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Martin:2015:GF

[Mar15i] P. A. Martin. The gamma function. In Higham et al. [HDG^{+}15], page 148. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Martin:2015:HL

[Mar15j] P. A. Martin. Hooke’s law. In Higham et al. [HDG^{+}15], pages 149–150. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Martin:2015:LE

[Mar15k] P. A. Martin. Laplace’s equation. In Higham et al. [HDG^{+}15], pages 155–156. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Martin:2015:SWE

[Mar15l] P. A. Martin. The shallowwater equations. In Higham et al. [HDG^{+}15], pages 167– 168. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Martin:2015:S

[Mar15m] P. A. Martin. Singularities. In Higham et al. [HDG^{+}15], pages 124–125. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Martin:2015:WE

[Mar15n] P. A. Martin. The wave equation. In Higham et al. [HDG^{+}15], page 171. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Martin:2015:WP

[Mar15o] P. A. Martin. Wave phenomena. In Higham et al. [HDG^{+}15], page 134. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Matalon:2015:FP

[Mat15] Moshe Matalon. Flame propagation. In Higham et al. [HDG^{+}15], pages 852–857. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Martin:2015:TM

[MB15] Dian I. Martin and Michael W.

Berry. Text mining. In Higham et al. [HDG^{+}15], pages 887– 891. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

McKee:2015:MPT

[McK15] Sean McKee. Modeling a pregnancy testing kit. In Higham et al. [HDG^{+}15], pages 864– 866. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

McPhedran:2015:EMT

[McP15] Ross C. McPhedran. Effective medium theories. In Higham et al. [HDG^{+}15], pages 500– 505. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Meini:2015:MC

[Mei15a] Beatrice Meini. Markov chains. In Higham et al. [HDG^{+}15], pages 116–117. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Meiss:2015:ODE

[Mei15b] James D. Meiss. Ordinary differential equations. In Higham et al. [HDG^{+}15], pages 181– 190. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Mendick:2015:MMR

[Men15] Heather Mendick. Mediated mathematics: Representations of mathematics in popular culture and why these matter. In Higham et al. [HDG^{+}15], pages 943– 952. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Miller:2015:PTA

[Mil15] Peter D. Miller. Perturbation theory and asymptotics. In Higham et al. [HDG^{+}15], pages 208– 218. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Moffatt:2015:FD

[Mof15a] H. K. Moffatt. Fluid dynamics. In Higham et al. [HDG^{+}15], pages 467–476. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Moffatt:2015:NSE

[Mof15b] H. K. Moffatt. The Navier– Stokes equations. In Higham et al. [HDG^{+}15], pages 162– 163. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Moro:2015:NA

[Mor15] Esteban Moro. Network analysis. In Higham et al. [HDG^{+}15], pages 360–374. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

McWhirter:2015:SP

[MP15] John G. McWhirter and Ian Proudler. Signal processing. In Higham et al. [HDG^{+}15], pages 533–545. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Marzouk:2015:UQ

[MW15] Youssef Marzouk and Karen Willcox. Uncertainty quantification. In Higham et al. [HDG^{+}15], pages 131–133. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Novick-Cohen:2015:CHE

[NC15] Amy Novick-Cohen. The Cahn– Hilliard equation. In Higham et al. [HDG^{+}15], pages 139– 140. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Nie:2015:SB

[Nie15a] Qing Nie. Systems biology. In Higham et al. [HDG^{+}15], pages 879–883. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Niederreiter:2015:RNG

[Nie15b] Harald Niederreiter. Random number generation. In Higham et al. [HDG^{+}15], pages 761– 762. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Oksendal:2015:ASA

[ØS15] Bernt Øksendal and Agnès Sulem. Applications of stochastic analysis. In Higham et al. [HDG^{+}15], pages 319–327. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Panfilov:2015:CM

[Pan15] Alexander V. Panfilov. Cardiac modeling. In Higham et al. [HDG^{+}15], pages 623–627. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Potter:2015:V

[PJ15] Kristin Potter and Chris R. Johnson. Visualization. In Higham et al. [HDG^{+}15], pages 843– 847. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Parolini:2015:S

[PQ15] Nicola Parolini and Alfio Quarteroni. Sport. In Higham et al. [HDG^{+}15], pages 598– 604. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Prosperetti:2015:B

[Pro15] Andrea Prosperetti. Bubbles. In Higham et al. [HDG^{+}15], pages 735–737. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Rantzer:2015:CT

[RÅ15] Anders Rantzer and Karl Johan Åström. Control theory. In Higham et al. [HDG^{+}15], pages 523–533. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Raydan:2015:NEN

[Ray15] Marcos Raydan. Nonlinear equations and Newton’s method. In Higham et al. [HDG^{+}15], pages 120–122. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Rockmore:2015:FFT

[Roc15] Daniel N. Rockmore. The fast Fourier transform. In Higham et al. [HDG^{+}15], pages 94–95. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL

http://press.princeton.edu/
titles/10592.html.

Saari:2015:BPA

[Saa15a] Donald G. Saari. From the N-body problem to astronomy and dark matter. In Higham et al. [HDG^{+}15], pages 771– 774. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Saari:2015:VS

[Saa15b] Donald G. Saari. Voting systems. In Higham et al. [HDG^{+}15], pages 891–895. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Santosa:2015:H

[San15a] Fadil Santosa. Homogenization. In Higham et al. [HDG^{+}15], page 103. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Santosa:2015:LSM

[San15b] Fadil Santosa. The level set method. In Higham et al. [HDG^{+}15], pages 114–116. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

                                        /press.princeton.edu/titles/
                                        10592.html.

Santosa:2015:MM

[San15c] Fadil Santosa. Multiscale modeling. In Higham et al. [HDG^{+}15], pages 119–120. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Sapiro:2015:MIP

[Sap15] Guillermo Sapiro. Mathematical image processing. In Higham et al. [HDG^{+}15], pages 813– 816. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Scheel:2015:PF

[Sch15] Arnd Scheel. Pattern formation. In Higham et al. [HDG^{+}15], pages 458–467. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Shuckburgh:2015:ESD

[Shu15] Emily Shuckburgh. Earth system dynamics. In Higham et al. [HDG^{+}15], pages 485–500. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Simoncini:2015:KS

[Sim15] Valeria Simoncini. Krylov subspaces. In Higham et al. [HDG^{+}15], pages 113–114. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Smith:2015:I

[Smi15] Malcolm C. Smith. Inerters. In Higham et al. [HDG^{+}15], pages 604–609. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Snieder:2015:IEU

[Sni15] Roel Snieder. Imaging the Earth using Green’s theorem. In Higham et al. [HDG^{+}15], pages 857– 860. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

                                        URL http://press.princeton.
                                        edu/titles/10592.html.

Sottile:2015:AG

[Sot15] Frank Sottile. Algebraic geometry. In Higham et al. [HDG^{+}15], pages 570–579. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Santosa:2015:IP

[SS15] Fadil Santosa and William W. Symes. Inverse problems. In Higham et al. [HDG^{+}15], pages 327–335. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Stewart:2015:HWG

[Ste15a] Ian Stewart. How to write a general interest mathematics book. In Higham et al. [HDG^{+}15], pages 906–912. ISBN 0-691-15039-7 (hardcover). LCCN QA155

                                        .P75 2015. URL http://press.
                                        princeton.edu/titles/10592.
                                        html.

Stewart:2015:SAM

[Ste15b] Ian Stewart. Symmetry in applied mathematics. In Higham et al. [HDG^{+}15], pages 402– 410. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Strang:2015:SFM

[Str15a] Gilbert Strang. A symmetric framework with many applications. In Higham et al. [HDG^{+}15], pages 661–665. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Strang:2015:WTH

[Str15b] Gilbert Strang. What to teach and how? In Higham et al. [HDG^{+}15], pages 938–940. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Suli:2015:NSP

[Sül15]  Endre Süli.  Numerical solu-

tion of partial differential equations. In Higham et al. [HDG^{+}15], pages 306–319. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Temme:2015:SF

[Tem15] Nico M. Temme. Special functions. In Higham et al. [HDG^{+}15], pages 227–235. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Tong:2015:CM

[Ton15] David Tong. Classical mechanics. In Higham et al. [HDG^{+}15], pages 374–383. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Tucker:2015:CAP

[Tuc15a] Warwick Tucker. Computeraided proofs via interval analysis. In Higham et al. [HDG^{+}15], pages 790–795. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Tucker:2015:IA

[Tuc15b] Warwick Tucker. Interval analysis. In Higham et al. [HDG^{+}15], pages 105–106. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Turner:2015:CMP

[Tur15] Peter R. Turner. Computation, modeling, and projects. In Higham et al. [HDG^{+}15], pages 935– 938. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Upton:2015:DFG

[Upt15] P. J. Upton. The delta function and generalized functions. In Higham et al. [HDG^{+}15], pages 139–141. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Verdu:2015:IT

[Ver15] Sergio Verdú. Information theory. In Higham et al. [HDG^{+}15], pages 545–552. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Villani:2015:KT

[VM15] Cédric Villani and Clément Mouhot. Kinetic theory. In Higham et al. [HDG^{+}15], pages 428–446. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Vygen:2015:CO

[Vyg15] Jens Vygen. Combinatorial optimization. In Higham et al. [HDG^{+}15], pages 564–570. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:/

/press.princeton.edu/titles/
10592.html.

Wampler:2015:R

[Wam15] Charles W. Wampler. Robotics.

In Higham et al. [HDG^{+}15], pages 767–769. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Wang:2015:IF

[Wan15] Z. Jane Wang. Insect flight. In Higham et al. [HDG^{+}15], pages 743–746. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Weaire:2015:F

[WH15] Denis Weaire and Stefan Hutzler. Foams. In Higham et al. [HDG^{+}15], pages 737– 741. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Willinger:2015:CN

[Wil15] Walter Willinger. Communication networks. In Higham et al. [HDG^{+}15], pages 883– 887. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Winkler:2015:ACG

[Win15] Peter Winkler. Applied combinatorics and graph theory. In Higham et al. [HDG^{+}15], pages 552–564. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Wright:2015:CON

[Wri15] Stephen J. Wright. Continuous optimization (nonlinear and linear programming). In Higham et al. [HDG^{+}15], pages 281– 293. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015.

URL http://press.princeton.
edu/titles/10592.html.

Yuan:2015:HCM

[xY15] Ya xiang Yuan. How Chinese mathematicians influence government. In Higham et al. [HDG^{+}15], pages 953–954. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.

Ying:2015:BPF

[Yin15] Lexing Ying. The N-body problem and the fast multipole method. In Higham et al. [HDG^{+}15], pages 775–778. ISBN 0-691-15039-7 (hardcover). LCCN QA155

.P75 2015. URL http://press.
princeton.edu/titles/10592.
html.